# Enhancing Security in Multi-Robot Systems through Co-Observation Planning, Reachability Analysis, and Network Flow

### v1

@@ II-B Co-observation schedule constraint
$\mathcal{Z}$ =$=\{\mathbf{z}\mid\lVert\mathbf{z}\rVert\leq d_{max}\},$ (13)
$\Pi_{\mathcal{Z}}(z)$ =$=\begin{cases}d_{max}\frac{\mathbf{z}}{\lVert\mathbf{z}\rVert}&\text{if }\lVert\mathbf{z}\rVert>d_{max},\\ \mathbf{z}&otherwise,\end{cases}$ (14)
where $a,b$ are the indices of the pair of agents required for a mutual inspection.
@@ II-C Definition of ellipsoidal reachability regions
In this section, we define ellipsoidal reachability regions based on pairs of co-observation locations on a trajectory, and a transformation of such region in axis-aligned form. Different types of reachability constraints (between an ellipsoid and different geometric entities such as a point, line, line segment, and polygon) alongside their projections, and differentials are introduced in subsequent sections. Together with the co-observation constraint 1, these will be used in the ADMM formulation of section II-A3 to provide security for all observed and unobservered periods.
@@
The reachability region for two waypoints $q(t_{1})=q_{1}$, $q(t_{2})=q_{2}$ is defined as the sets of points $q^{\prime}$ in the workspace such that there exist a trajectory $q(t)$ where $q(t^{\prime})=q^{\prime}$, $t_{1}\leq t^{\prime}\leq t_{2}$ and $q(t)$ satisfies the velocity constraint $d(q(t),q(t+1))\leq v_{max}$.
This region can be analytically bounded via an ellipsoid:
@@
The reachability ellipsoid $\mathcal{E}$ is defined as the region $\mathcal{E}(q_{1},q_{2},t_{1},t_{2})=\{\tilde{q}\in\mathbb{R}^{n}:d(q_{1},\tilde{q})+d(\tilde{q}+q_{2})<2a\}$, where $a=\frac{v_{maq}}{2}(t_{2}-t_{1})$.
@@
Figure 2: Figure 2a showcases the ellipsodial reachability region. Black line is the planned trajectory, $q_{1}$ and $q_{2}$ are two co-observed locations the robot are expected at given time $t_{1}$ and $t_{2}$, dashed lines show possible trajectory of a compromised robot during unobserved period. Axis of global frame $\mathcal{F}$ and canonical frame $\mathcal{F}_{\mathcal{E}}$ are shown as $x_{\mathcal{F}},y_{\mathcal{F}}$ and $x_{\mathcal{E}},y_{\mathcal{E}}$ respectively. In Figure 2b, a point $q_{avoid}$ inside ellipsoid is projected to the areas outside the ellipsoid $f_{p}$. In Figure 2c, the projection is simplified to the point-ellipse constraint as projecting a point inside ellipse $p_{L}$ to $p_{t}$ on outside the ellipse. In Figure 2d, this constraint is simplified to either a plane-ellipse constraint $D_{1}$ or a point-ellipse constraint $D_{2}$.
The region $\mathcal{E}(q_{1},q_{2})$ is an ellipsoid with foci at $q_{1},q_{2}$, center $o_{\mathcal{E}}=\frac{1}{2}(q_{1}+q_{2})$, and the major radius equal to $a$. We denote as $c_{\mathcal{E}}=\frac{1}{2}\lVert q_{1}-q_{2}\rVert=\lVert o_{\mathcal{E}}-q_{1}\rVert$ the semi-axis distance from the center to a foci. This reachability ellipsoid is an over-approximation of the exact reachability region in Definition 3.
Expanding upon the concept of co-observation and reachability, Definition 1 can be written as,
Remark 1
A multi-robot trajectory is secured against plan-deviation attacks if there exist a co-observation plan such that the reachability region between each consecutive co-observation does not intersect with any forbidden regions.
@@ II-D Transformation to canonical coordinates
To simplify the problem, we employ a canonical rigid body transformation that repositions the ellipse $\mathcal{E}$ from the global frame $\mathcal{F}$ to a canonical frame $\mathcal{F}_{\mathcal{E}}$, and perform this transformation in a differentiable manner with respect to the two foci serving as waypoints for a robot. The canonical frame $\mathcal{F}_{\mathcal{E}}$ is defined with the origin at the ellipsoid’s center, and the first axis aligned with the foci Figure 2a.
For all reachability ellipsoid constraints, we first solve the problem in the canonical frame $\mathcal{F}_{\mathcal{E}}$, and transform the solutions to the global frame $\mathcal{F}$ using the transformation (15). We parametrize the transformation from $\mathcal{F}$ to $\mathcal{F}_{\mathcal{E}}$ using a rotation $R^{\mathcal{F}}_{\mathcal{E}}$ and a translation $o^{\mathcal{F}}_{\mathcal{E}}$, which, to simplify the notation, we refer to as $R$ and $o$, respectively. The transformation of a point from the frame $\mathcal{F}_{\mathcal{E}}$ to the frame $\mathcal{F}$ and its inverse are given by the rigid body transformations:
${q^{\mathcal{F}}}=R{q^{\mathcal{E}}}+o,\quad{q^{\mathcal{E}}}=R^{\mathrm{T}}({q^{\mathcal{F}}}-o).$ (15)
We define $\nu_{\mathcal{F}}$ and $\nu_{\mathcal{E}}$ to represent the $x$-axis unitary vector of $\mathcal{F}_{\mathcal{E}}$ in the frames $\mathcal{F}$ and $\mathcal{F}_{\mathcal{E}}$, respectively. Formally:
$\nu^{\prime}_{\mathcal{F}}$ $=q_{2}-q_{1},$ $\nu_{\mathcal{F}}$ $=\frac{\nu^{\prime}_{\mathcal{F}}}{\lVert\nu_{\mathcal{F}}\rVert},$ $\nu_{\mathcal{E}}$ $=[1,0,0]^{T};$ (16)
see Fig.2a for an illustration. Note that $\nu_{\mathcal{E}}$ is constant while, for the sake of clarity, $\nu_{\mathcal{F}}$ depends on $q_{1},q_{2}$. From the definition of $\mathcal{F}_{\mathcal{E}}$, the rotation $R$ can be found using a Householder rotation, while from (15) the translation $o$ must be equal to the center of $\mathcal{E}(q_{1},q_{2})$ expressed in $\mathcal{F}$, i.e.:
$R$ $=H(\nu_{\mathcal{F}}(q_{1},q_{2}),\nu_{\mathcal{E}}(q_{1},q_{2})),$ $o$ $=\frac{1}{2}(q_{1}+q_{2}).$ (17)
To simplify the notation, in the following, we will consider $H$ to be a function of $q_{1},q_{2}$ directly, i.e. $H(q_{1},q_{2})$. The ellipse $\mathcal{E}$ expressed in $\mathcal{F}_{\mathcal{E}}$ is given by =${\mathcal{E}^{\mathcal{E}}}=\{{q^{\mathcal{E}}}\in m:d(q^{\mathcal{E}}_{1},{q^{\mathcal{E}}})+d({q^{\mathcal{E}}},q^{\mathcal{E}}_{2})<2a\}$, with foci $q^{\mathcal{E}}_{1},q^{\mathcal{E}}_{2}$ in $\mathcal{F}_{\mathcal{E}}$ defined as
$q^{\mathcal{E}}_{1}=\begin{bmatrix}c&0&0\end{bmatrix}^{\mathrm{T}},$ $q^{\mathcal{E}}_{2}=\begin{bmatrix}-c&0&0\end{bmatrix}^{\mathrm{T}},$ (20)
and semi-axis distance $c=\frac{\lVert q_{2}-q_{1}\rVert}{2}$.
Definition 5
The reachability ellipsoid $\mathcal{E}$ in the canonical frame is defined as as the zero level set of the quadratic function
where
$Q=\operatorname{diag}(a^{-2},b^{-2},b^{-2}),$ (22)
Lemma 1
${E^{\mathcal{F}}}({q^{\mathcal{F}}})=({q^{\mathcal{F}}}-o_{\mathcal{E}})^{\mathrm{T}}H^{\mathrm{T}}QH({q^{\mathcal{F}}}-o_{\mathcal{E}})-1$ (23)
Proof:
We are now ready to formulate constraints based on reachability ellipsoids against different types of forbidden regions: a point, a plane, a segment, and a convex polygon which is is done by defining $D(q)$, its differential $\partial_{q}D(q)$, the set $\zeta$ and the projection $\Pi_{\zeta}$.
@@ II-E Point-ellipsoid reachability constraint
As shown in Fig.2b, we consider a forbidden region in the shape of a single point $q_{avoid}$. The goal is to design the trajectory $q(t)$ such that $q_{avoid}\notin\mathcal{E}(q_{1},q_{2})$. This constraint can be written as,
@@
[…]
$\mathcal{Z}$ =$=\{q\in\mathbb{R}^{nm}:\lVert D_{p}(\mathbf{q})\rVert=0\},$ (25)
@@
$\pi_{p\mathcal{E}}=\operatorname*{\arg\!\min}_{q_{p}\in\mathcal{E}^{c}}\quad\lVert q_{avoid}-q_{p}\rVert^{2}$ (27)
@@
$q^{\mathcal{E}}_{p}={(I+sQ)}^{-1}q^{\mathcal{E}}_{avoid}=Sq^{\mathcal{E}}_{avoid}$ (28)
@@
${q^{\mathcal{E}}_{p}}^{\mathrm{T}}Qq^{\mathcal{E}}_{p}-1={q^{\mathcal{E}}_{avoid}}^{\mathrm{T}}Q^{\prime}(s)q^{\mathcal{E}}_{avoid}-1=0$ (29)
@@
$Q^{\prime}(s)=S^{\mathrm{T}}QS=\operatorname{diag}\left(\frac{a^{2}}{(s+a^{2})^{2}},\frac{b^{2}}{(s+b^{2})^{2}},\frac{b^{2}}{(s+b^{2})^{2}}\right)$ (30)
@@
$\pi_{p\mathcal{E}}(q)=R^{-1}(q(t_{1}),q(t_{2}))q^{\mathcal{E}}_{p}+o\\ =R^{-1}(q(t_{1}),q(t_{2}))Sq^{\mathcal{E}}_{avoid}+o\\ =R^{-1}SR(q_{avoid}-o)+o$ (31)
@@
$\partial_{\left[\begin{smallmatrix}q_{1}\\ q_{2}\end{smallmatrix}\right]}\pi_{p\mathcal{E}}=-2H[SH(q-o)]_{\times}U\\ +\left((q^{\mathrm{T}}\partial_{s}Q^{\prime}q)^{-1}H^{-1}Q^{\prime}qq^{\mathrm{T}}(4Q^{\prime}H[q-o]_{\times}U\right.\\ \left.+2Q^{\prime}H\partial_{q}o-\partial_{b}Q^{\prime}qq\partial_{q}b)-sH^{-1}S^{2}\partial_{b}Qq\partial_{q}b\right)\\ -2H^{-1}SH[q-o]_{\times}U+(H^{-1}SH-I)\partial_{q}o$ (32)
@@ II-F Plane-ellipsoid reachability constraint
For forbidden region in the shape of a hyperplane $\mathcal{L}({q^{\mathcal{F}}})=\{{q^{\mathcal{F}}}\in\mathbb{R}^{n}:\mathbf{n}^{\mathrm{T}}{q^{\mathcal{F}}}=d\}$ (as shown in Figure 2c), the reachability constraint can be defined as $\mathcal{L}\cap\mathcal{E}(q_{1},q_{2},a)=\emptyset$. When transformed into the canonical frame, the hyperplane can be written as ${\mathcal{L}^{\mathcal{E}}}({q^{\mathcal{E}}})=\{{q^{\mathcal{E}}}\in\mathbb{R}^{m}:\mathbf{n}_{\mathcal{E}}^{\mathrm{T}}{q^{\mathcal{E}}}=d_{\mathcal{E}}\}$, where,
$\mathbf{n}_{\mathcal{E}}=H(q_{1},q_{2})\mathbf{n},\quad d_{\mathcal{E}}=-\mathbf{n}^{\mathrm{T}}o+d.$ (33)
Intuitively, $d_{\mathcal{E}}$ can be thought as a distance between the the plane $\mathcal{L}$ and the origin (i.e., the center of $\mathcal{E}$).
For every ${\mathcal{L}^{\mathcal{E}}}({q^{\mathcal{E}}})$, there always exist two planes that are both parallel to $\mathcal{L}$ and tangential to the ellipse (i.e. resulting in a unique intersection point), $\mathcal{L}_{1}^{\mathcal{E}}=\{{q^{\mathcal{E}}}\in\mathbb{R}^{m}:\mathbf{n}_{\mathcal{E}}^{\mathrm{T}}{q^{\mathcal{E}}}=d_{\mathcal{E}t}\}$ and $\mathcal{L}_{2}^{\mathcal{E}}=\{{q^{\mathcal{E}}}\in\mathbb{R}^{m}:\mathbf{n}_{\mathcal{E}}^{\mathrm{T}}{q^{\mathcal{E}}}=-d_{\mathcal{E}t}\}$ (Figure 2c). And the intersection point can be written as:
@@
where $d_{\mathcal{E}t}=\sqrt{n_{\mathcal{E}}^{\mathrm{T}}Q^{-1}n_{\mathcal{E}}}$; intuitively, $d_{\mathcal{E}t}$ can be thought as a distance between the tangent plane $\mathcal{L}_{1}^{\mathcal{E}}$ (or $\mathcal{L}_{2}^{\mathcal{E}}$) and the origin (i.e., the center of the ellipse $\mathcal{E}$).
We introduce the concept of a tangent interpolation point on the hyperplane to characterize the relationship between the plane and the ellipse.
@@
The tangent interpolation point $p^{\mathcal{E}}_{\mathcal{L}}\in\mathcal{L}$ is defined on the plane by
@@
Intuitively, the point $p^{\mathcal{E}}_{\mathcal{L}}$ is the point on $\mathcal{L}$ which is closest to either $p^{\mathcal{E}}_{1}$ or $p^{\mathcal{E}}_{2}$. Note that when $d_{\mathcal{E}}=d_{\mathcal{E}t}$ or $d_{\mathcal{E}}=-d_{\mathcal{E}t}$, $p^{\mathcal{E}}_{\mathcal{L}}=p^{\mathcal{E}}_{t_{1}}$ or $p^{\mathcal{E}}_{\mathcal{L}}=p^{\mathcal{E}}_{t_{2}}$, respectively. When $d_{\mathcal{E}}\in[-d_{\mathcal{E}t},d_{\mathcal{E}t}]$, the plane $\mathcal{L}$ and the ellipsoid $\mathcal{E}$ have at least one intersection, thus violating our desired reachability constraint.
@@
$\mathcal{Z}$ =$=\{\mathbf{q}\in\mathbb{R}^{nm}:\lVert D_{\mathbf{n}}(\mathbf{q})\rVert=0\},$ (37)
@@
Based on the Proposition 4, the differential of (36) can be written as:
$\partial_{q}{D_{\mathbf{n}\mathcal{E}}}=-2H[{\Pi^{\mathcal{E}}}_{\mathbf{n}\mathcal{E}}]_{\times}\!M+H^{-1}\partial_{q}{\Pi^{\mathcal{E}}}_{\mathbf{n}\mathcal{E}}$ (43)
II-G Line-segment-ellipse reachability constraint
For a intermediate step to get the relationship between a polygon shaped forbidden region, the relative position between the ellipse and segment of the hyperplane that defines the region is studied. Assume the segment on hyperplane ${\mathcal{L}^{\mathcal{E}}}({q^{\mathcal{E}}})=\{{q^{\mathcal{E}}}\in\mathbb{R}^{m}:\mathbf{n}_{\mathcal{E}}^{\mathrm{T}}{q^{\mathcal{E}}}=d_{\mathcal{E}}\}$ has endpoints ${p_{1}^{\mathcal{E}}}$ and ${p_{2}^{\mathcal{E}}}$; then the segment can be defined as:

### v2

@@ II-B Co-observation schedule constraint
$\mathcal{Z}$ $=\{\mathbf{z}\mid\lVert\mathbf{z}\rVert\leq d_{max}\},$ (8)
$\Pi_{\mathcal{Z}}(z)$ $=\begin{cases}d_{max}\frac{\mathbf{z}}{\lVert\mathbf{z}\rVert}&\text{if }\lVert\mathbf{z}\rVert>d_{max},\\ \mathbf{z}&otherwise,\end{cases}$ (9)
where $a,b$ are the indices of the pair of agents required for a mutual inspection, $d_{max}$ is the maximum distance between a pair of robots that the security can be ensured through co-observation.
@@ II-C Definition of ellipsoidal reachability regions
In this section, we define ellipsoidal reachability regions with respect to a pair of locations along a trajectory. We then use the ellipsoids to introduce various reachability constraints with respect to other geometric entities, such as points, lines, line segments, and polygons. These constraints, combined with the co-observation constraint 1, form the basis of the secured ADMM planning is presented in Section II-A2.
@@
Consider a robot $i$ starting from $q_{1}$ at time $t_{1}$ and reaching $q_{2}$ at time $t_{2}$. The reachability region between $t_{1}$ and $t_{2}$ is defined as the set of points $q^{\prime}$ in the free configuration space for which there exists a kinematically feasible trajectory containing $q^{\prime}$.
For simplicity, in this paper we consider only a maximum velocity constraint $v_{max}$; the reachability region for $q^{\prime}$ can then be over-approximated by the following (the over-approximation is obtained by neglecting physical obstacles):·
@@
The reachability ellipsoid $\mathcal{E}$ is defined as the region $\mathcal{E}(q_{1},q_{2},t_{1},t_{2})=\{\tilde{q}\in\mathbb{R}^{n}:d(q_{1},\tilde{q})+d(\tilde{q},q_{2})<2a\}$, where $a=\frac{v_{max}}{2}(t_{2}-t_{1})$, and $d(\cdot,\cdot)$ denotes the Euclidean distance between two points.
This region is an ellipsoid because it represents the set of points $q^{\prime}$ whose sum of distances two foci $q_{1}$ and $q_{2}$ is less than $2a$, where $a$ is the major radius of the ellipsoid.
@@
Figure 3: (3a) The black line is the planned trajectory, $q_{1}$ and $q_{2}$ are two co-observed locations where the robots are expected at given times $t_{1}$ and $t_{2}$, dashed lines show a possible trajectory of a compromised robot during the unobserved period. The axis of global frame $\mathcal{F}$ and canonical frame $\mathcal{F}_{\mathcal{E}}$ are shown as $x_{\mathcal{F}},y_{\mathcal{F}}$ and $x_{\mathcal{E}},y_{\mathcal{E}}$ respectively. (3b) For point-ellipsoid constraint, $q_{avoid}$ is projected to the areas outside the ellipsoid to $q_{p}$. (3c) For plane-ellipsoid constraint, the projection is simplified to the point-ellipse constraint that projects point $p_{L}$ outside the ellipse to $p_{t}$. (3d) For convex-polygon-ellipsoid constraint, the projection is either a plane-ellipse constraint $D_{1}$ (for the red region) or a point-ellipse constraint $D_{2}$ (for the brown region).
The condition that a reachability ellipsoid $\mathcal{E}(q_{1},q_{2})$ does not intersect with any forbidden region is sufficient to secure the trajectory of the robot between $q_{1}$ and $q_{2}$ according to Definition 1: any deviation to a point $p_{o}$ outside $\mathcal{E}(q_{1},q_{2})$ will cause the robot to miss a potential observation at $p_{2}$, $t_{2}$. Definition 1 can then be restated as follows:
Definition 4
A multi-robot trajectory is secured against plan-deviation attacks if there exists a co-observation plan such that the reachability region between each consecutive co-observation does not intersect with any forbidden regions.
@@ II-D Transformation to canonical coordinates
To use the definition of reachability ellipsoid from the section above as computational constraints for ADMM, we first apply a differentiable rigid body transformation to reposition the ellipse $\mathcal{E}$ from the global frame $\mathcal{F}$ to a canonical frame $\mathcal{F}_{\mathcal{E}}$, where the origin of $\mathcal{F}_{\mathcal{E}}$ is at the ellipsoid’s center $o=\frac{1}{2}(q_{1}+q_{2})$, and the first axis of $\mathcal{F}_{\mathcal{E}}$ is aligned with the foci (see Fig. 3a for an illustration). From this definition, a unit vector along the first axis of the ellipsoid in the frames $\mathcal{F}$ and $\mathcal{F}_{\mathcal{E}}$ is given by $\nu_{\mathcal{F}}=\frac{q_{2}-q_{1}}{\lVert q_{2}-q_{1}\rVert}$ and $\nu_{\mathcal{E}}=[1,0,0]^{T}$, respectively. The full transformation from coordinates $q^{\mathcal{E}}\in\mathcal{F}_{\mathcal{E}}$ to $q\in\mathcal{F}$, and its inverse, are then parametrized by a rotation $R^{\mathcal{F}}_{\mathcal{E}}$ and a translation $o^{\mathcal{F}}_{\mathcal{E}}$ as (we drop the subscript and superscript from $R^{\mathcal{F}}_{\mathcal{E}}$ and $o^{\mathcal{F}}_{\mathcal{E}}$ to simplify the notation):
$q=R{q^{\mathcal{E}}}+o,\quad{q^{\mathcal{E}}}=R^{\mathrm{T}}(q-o).$ (10)
where the rotation matrix $R=H(\nu_{\mathcal{F}}(q_{1},q_{2}),\nu_{\mathcal{E}})$ is a Householder rotation $H$ (a differentiable linear transformation describing the minimal rotation between the two vectors, see Section -B for details). To simplify the notation, in the following, we will consider $H$ to be a function of $q_{1},q_{2}$ directly, i.e. $H(q_{1},q_{2})$. More details on (10) are given in Section -A.
Reachability constraints are formulated with respect to different types of forbidden regions a point, a plane, a segment, and a convex polygon.
@@ II-E Point-ellipsoid reachability constraint
For a forbidden region in the shape of a single point $q_{avoid}$ (Fig. 3b), the constraint is written as:
@@
$D(\mathbf{q})$ $=\begin{cases}\pi_{p\mathcal{E}}(\mathbf{q})-q_{avoid}&q_{avoid}\in\mathcal{E},\\ {0}&\textrm{otherwise}.\end{cases}$ (11)
$\mathcal{Z}$ $=\{\mathbf{q}\in\mathbb{R}^{nm}:\lVert D_{p}(\mathbf{q})\rVert=0\},$ (12)
@@
$\pi_{p\mathcal{E}}=\operatorname*{\arg\!\min}_{q_{p}\in\mathcal{E}^{c}}\quad\lVert q_{avoid}-q_{p}\rVert^{2},$ (14)
@@
$q^{\mathcal{E}}_{p}={(I+sQ)}^{-1}q^{\mathcal{E}}_{avoid}=Sq^{\mathcal{E}}_{avoid},$ (15)
@@
${q^{\mathcal{E}}_{p}}^{\mathrm{T}}Qq^{\mathcal{E}}_{p}-1={q^{\mathcal{E}}_{avoid}}^{\mathrm{T}}Q^{\prime}(s)q^{\mathcal{E}}_{avoid}-1=0,$ (16)
@@
$Q^{\prime}(s)=S^{\mathrm{T}}QS=\operatorname{diag}\left(\frac{a^{2}}{(s+a^{2})^{2}},\frac{b^{2}}{(s+b^{2})^{2}},\frac{b^{2}}{(s+b^{2})^{2}}\right).$ (17)
@@
$\pi_{p\mathcal{E}}(q)=R^{-1}(q(t_{1}),q(t_{2}))q^{\mathcal{E}}_{p}+o\\ =R^{-1}(q(t_{1}),q(t_{2}))Sq^{\mathcal{E}}_{avoid}+o\\ =R^{-1}SR(q_{avoid}-o)+o.$ (18)
@@
$\partial_{\left[\begin{smallmatrix}q_{1}\\ q_{2}\end{smallmatrix}\right]}\pi_{p\mathcal{E}}=-2H[SH(q-o)]_{\times}U\\ +\left((q^{\mathrm{T}}\partial_{s}Q^{\prime}q)^{-1}H^{-1}Q^{\prime}qq^{\mathrm{T}}(4Q^{\prime}H[q-o]_{\times}U\right.\\ \left.+2Q^{\prime}H\partial_{q}o-\partial_{b}Q^{\prime}qq\partial_{q}b)-sH^{-1}S^{2}\partial_{b}Qq\partial_{q}b\right)\\ -2H^{-1}SH[q-o]_{\times}U+(H^{-1}SH-I)\partial_{q}o.$ (19)
@@ II-F Plane-ellipsoid reachability constraint
For a hyperplane shaped forbidden region $\mathcal{L}(q)=\{q\in\mathbb{R}^{m}:\mathbf{n}^{\mathrm{T}}q=d\}$ (Fig. 3c), the reachability constraint is $\mathcal{L}\cap\mathcal{E}(q_{1},q_{2},a)=\emptyset$. When transformed into the canonical frame, the hyperplane can be written as ${\mathcal{L}^{\mathcal{E}}}({q^{\mathcal{E}}})=\{{q^{\mathcal{E}}}\in\mathbb{R}^{m}:\mathbf{n}_{\mathcal{E}}^{\mathrm{T}}{q^{\mathcal{E}}}=d_{\mathcal{E}}\}$, with $\mathbf{n}_{\mathcal{E}}=H(q_{1},q_{2})\mathbf{n},\quad d_{\mathcal{E}}=-\mathbf{n}^{\mathrm{T}}o+d$.
For every ${\mathcal{L}^{\mathcal{E}}}({q^{\mathcal{E}}})$, there exist two planes that are both parallel to $\mathcal{L}$ and tangential to the ellipse (i.e. resulting in a unique intersection point Fig. 3c), $\mathcal{L}_{1}^{\mathcal{E}}=\{{q^{\mathcal{E}}}\in\mathbb{R}^{m}:\mathbf{n}_{\mathcal{E}}^{\mathrm{T}}{q^{\mathcal{E}}}=d_{\mathcal{E}t}\}$ and $\mathcal{L}_{2}^{\mathcal{E}}=\{{q^{\mathcal{E}}}\in\mathbb{R}^{m}:\mathbf{n}_{\mathcal{E}}^{\mathrm{T}}{q^{\mathcal{E}}}=-d_{\mathcal{E}t}\}$. The intersection point can be written as:
@@
where $d_{\mathcal{E}t}=\sqrt{n_{\mathcal{E}}^{\mathrm{T}}Q^{-1}n_{\mathcal{E}}}$; intuitively, $d_{\mathcal{E}t}$ can be treated as a distance between the tangent plane $\mathcal{L}_{1}^{\mathcal{E}}$ (or $\mathcal{L}_{2}^{\mathcal{E}}$) and the origin (i.e., the center of the ellipse $\mathcal{E}$). The concept of tangent interpolation points are introduced to characterize the relationship between the plane and the ellipsoid.
@@
The tangent interpolation point $p^{\mathcal{E}}_{\mathcal{L}}\in\mathcal{L}$ between the plane $\mathcal{L}$ and the ellipsoid $\mathcal{E}$ is defined on the plane by
@@
Intuitively, the point $p^{\mathcal{E}}_{\mathcal{L}}$ is the closest point on $\mathcal{L}$ to either $p^{\mathcal{E}}_{1}$ or $p^{\mathcal{E}}_{2}$. Note that when $d_{\mathcal{E}}=d_{\mathcal{E}t}$ or $d_{\mathcal{E}}=-d_{\mathcal{E}t}$, $p^{\mathcal{E}}_{\mathcal{L}}=p^{\mathcal{E}}_{t_{1}}$ or $p^{\mathcal{E}}_{\mathcal{L}}=p^{\mathcal{E}}_{t_{2}}$, respectively. When $d_{\mathcal{E}}\in[-d_{\mathcal{E}t},d_{\mathcal{E}t}]$, the plane $\mathcal{L}$ and the ellipsoid $\mathcal{E}$ have at least one intersection, thus violating our desired reachability constraint.
@@
$\mathcal{Z}$ $=\{\mathbf{q}\in\mathbb{R}^{nm}:\lVert D_{\mathbf{n}}(\mathbf{q})\rVert=0\},$ (23)
@@
The differential of (22) can be written as:
$\partial_{q}{D_{\mathbf{n}\mathcal{E}}}=-2H[{\Pi^{\mathcal{E}}}_{\mathbf{n}\mathcal{E}}]_{\times}\!M+H^{-1}\partial_{q}{\Pi^{\mathcal{E}}}_{\mathbf{n}\mathcal{E}}.$ (29)
II-G Convex-polygon-ellipse reachability constraint
The reachability constraint for a convex polygon is treated as a union of reachability constraints for each individual segments that define the hyperplane. We define
ADMM constraint 4 (Convex-polygon-ellipsoid reachability constraint)
$\mathcal{Z}$ $=\{\mathbf{q}\in\mathbb{R}^{nm}:\lVert D(\mathbf{q})\rVert=0\},$ (31)
$\Pi_{\mathcal{Z}}(\mathbf{z})$ $=0,$ (32)
where $D_{seg}$ are the constraint functions for all line segments that define the convex polygon region. For hyperplane ${\mathcal{L}^{\mathcal{E}}}({q^{\mathcal{E}}})=\{{q^{\mathcal{E}}}\in\mathbb{R}^{m}:\mathbf{n}_{\mathcal{E}}^{\mathrm{T}}{q^{\mathcal{E}}}=d_{\mathcal{E}}\}$ with endpoints $p_{1}^{\mathcal{E}}$} and $p_{2}^{\mathcal{E}}$}; this segment is defined as: